# Modeling Sustainable City Trips: Integrating CO2e Emissions, Popularity, and Seasonality into Tourism Recommender Systems

### v2

@@ 5.2. Estimating Seasonal Demand
In the context of seasonal tourism demand, studies suggest that the average room price is one of the pivotal business performance indicators in the hotel industry (Israeli, 2002; O’Neill and Carlbäck, 2011; Pine and Phillips, 2005). We calculate the seasonality Gini index for a city using Gini coefficients derived from non-financial (NFI) and financial indicators (FI), as described by Suštar and Ažić (2020). Non-financial indicators consist of monthly counts of arriving visitors (AVC) from TourMIS, while financial indicators involve the ADR computed from Airbnb listings. The Gini coefficient values for each indicator span from zero to one, with zero signifying a complete lack of seasonality and seems to be an active or equal distribution of volumes all year round. A Gini coefficient of one indicates complete seasonality, i.e., the total volume is registered in one single month (Suštar and Ažić, 2020). Gini coefficients are computed annually for AVC, while monthly calculations are performed for ADR.
@@
(5) $\sigma(c_{i}^{j})=\Gamma_{AVC}\cdot G_{AVC}^{(i)}+\Gamma_{ADR}\cdot G_{ADR}^{(i)(j)}$
@@
Table 4. Analysis of the relationship between Airbnb ADR and TourMIS AVC: Pearson’s correlation coefficients and T-Test results for selected European cities
@@ Results
We calculate the Gini coefficients for the AVC and ADR, presenting the results in Table 5. Madrid exhibits the least seasonality in AVC, suggesting consistent demand throughout the year, while Munich registers the highest seasonality. Additionally, Brussels displays minimal seasonality in ADR (close to 0), and Munich exhibits maximum ADR seasonality in September and October. The heightened seasonality in Munich’s AVC and ADR can also be attributed to the annual Oktoberfest event occurring in September (Herrmann and Herrmann, 2014).
Table 5. Gini coefficients illustrating the seasonality indicators in selected cities. The bold notations highlight the maximum Gini indices for ADR in the city, while the italicized and underlined ones denote the minimum
@@ 5.3. Summary
According to research, the most effective way to measure tourism seasonality is by computing the Gini coefficients of various financial and non-financial indicators (Suštar and Ažić, 2020). We collected data from TourMIS and Airbnb to analyze non-financial indicators such as monthly arrival visitor counts (AVC) and financial indicators like average daily rates (ADR), respectively. The seasonal demand index is a weighted combination of Gini coefficients calculated from non-financial and financial indicators.
@@
6. User Perception of Sustainable City Trips
To explore how users perceive sustainability when looking for city trip recommendations, we conducted a user study involving participants with diverse travel experiences and preferences. This method provides valuable insights into decision-making intricacies, as established by prior research (Wilson, 1981). By directly engaging real-world participants in simulated scenarios, we aimed to discern how tourists assess different aspects of a city trip, negotiate trade-offs, align preferences with sustainable tourism practices, and assign weights to criteria in their decision-making.
6.1. User Study Design
@@
The questionnaire was designed using Qualtrics Experience Management Software  [^13] https://www.qualtrics.com, an online survey platform. We recruited participants through the online crowdsourcing platform Prolific  [^14] https://www.prolific.com, renowned for its efficacy in subject recruitment for the scientific community (Palan and Schitter, 2018). With a focus on European participants who listed travel as one of their hobbies, the questionnaire, designed in English, was distributed to individuals through Prolific’s advanced pre-screening options, and 200 final responses were collected. To ensure gender diversity, the preset distribution aimed for an equal representation of 50% males and 50% females. Demographic analyses of the survey data indicated that 33.8% of the participants fell within the 25-34 age group, followed by 24.3% in the 18-24 age group, and the remaining were above the age of 35.
6.2. Transportation Sustainability Concerns
@@
Figure 4. CO2e, travel time (TT), estimated distance, and cost as shown to the users for each trip scenario, separated by transportation mode. Each city in every trip scenario is denoted by its respective IATA code
@@
Figure 5. Distribution of user responses (%) for different transport modes in various trip scenarios. Each city in every trip scenario is denoted by its respective IATA code
We investigated the reasons behind selecting trade-offs for various transportation modes in each trip scenario, as illustrated in Figure 6. Our analysis reveals valuable insights from survey responses. For train travel, a significant 47.1% of respondents prioritize convenience, while 15.65% emphasize selecting the mode with the least CO2e emissions. Affordability is a critical factor for 13.91% of participants, and 8.70% opt for the cheapest option available. Additionally, 5.22% favor trains over flights for cost considerations. In contrast, driving is primarily chosen for its convenience and flexibility, with an overwhelming 73.81% of respondents highlighting this aspect. Affordability remains a factor, as 18.57% opt for the cheapest driving option. Some respondents (6.67%) perceive driving as environmentally better than flying. Flying is chosen by 76.53% for its speed, while 11.91% remains unconcerned about CO2e emissions. The findings underline the multifaceted nature of decision-making, encompassing convenience, environmental concerns, and cost considerations across different transportation modes. Our results are consistent with those presented by Avogadro et al. (2021), which also indicates a preference for public transport over flying for shorter distances.
Figure 6. Reasonings provided by respondents when choosing modes of transport across all trips
6.3. Understanding Trade-offs
The user study also aimed to explore the various trade-offs associated with trip planning. Participants were presented with Likert scale (Joshi et al., 2015) statements from ”not at all important” to ”extremely important” to gauge their agreement levels to the following statements:
@@ 7. $S_{7}$:
Figure 7. Degree of the importance of the various factors influencing decision-making during trip planning
The analysis of user responses, as depicted in Figure 7, provides insights into the trade-offs users are willing to make when considering various attributes in travel decision-making. Notably, for the attribute ”popular but crowded,” a minimal percentage (2.99%) strongly agreed, while the majority (31.34%) agreed. Conversely, for the attribute ”attractive but popular”, a significant proportion (55.22%) agreed, indicating a higher tolerance for popularity in the pursuit of attractiveness. The consideration of ”overall attractiveness” saw 51.24% in agreement. Respondents expressed varying opinions on the ”visiting in peak season”, with 1.99% strongly disagreeing and 33.83% expressing disagreement. Furthermore, factors such as ”off-season cost savings”, ”Climate at the destination” during the time of travel, and ”off-season discounts” revealed nuanced preferences, with notable percentages in agreement (51.74%, 37.31%, 39.30%, respectively) and distinct proportions holding dissenting views. These results contribute valuable insights into the varied considerations influencing users’ travel preferences.
7. Societal Fairness
This section explores the concept of Societal Fairness (S-Fairness), which is a crucial component in evaluating city trip recommendations. S-Fairness seeks to ensure that tourism benefits and impacts are distributed fairly among a broad range of stakeholders, including not only tourists, service providers, and platforms, but also non-participating entities such as residents, locals, and the environment (Banerjee, 2023; Banik et al., 2023).
@@ 7. Societal Fairness
7.1. Defining the S-Fairness Indicator
@@
(6) $SFI(c_{i}^{j})=\alpha\cdot Z(c_{i})+\beta\cdot\rho(c_{i})+\Gamma\cdot\sigma(c_{i}^{j})$
@@
7.2. Learning Weights of the Indices
One of the main goals of the user study described in Section 6.1 was to understand the importance of different factors that the users consider when choosing travel destinations. Participants were requested to prioritize their transportation choices based on travel time, CO2e footprint, and cost. They also provided feedback on other essential factors, such as the number of points of interest, Google image search values, and the total number of Tripadvisor reviews and opinions. Additionally, they indicated the importance of costs associated with available accommodation and crowd levels in the city during their visit. These factors align with the coefficients of the ”emissions trade-off index” $\mathit{Z(c_{i})}$ in Equation 1, ”popularity index” $\rho(c_{i})$ in Equation 3, and ”seasonality index” $\sigma(c_{i}^{j})$ in Equation 5 respectively.
@@ 7.2. Learning Weights of the Indices
We calculated weighted averages on Likert scales, spanning from ”not important at all”, having a minimum weight of 1, to ”extremely important”, with a maximum weight of 5, enabling us to identify patterns in the composite indicator. The distribution of the absolute values of these weights, obtained through the weighted average of Likert scale results, are shown in  Figure 8. Additionally, we applied min-max normalization (Patro and Sahu, 2015) to normalize these averages within each category, enabling us to gauge their relative significance in their respective categories. These normalized weights offer valuable insights into participants’ preferences and priorities, shedding light on the factors that significantly influence their decision-making when choosing travel destinations. After incorporating the normalized weights, the expressions for the emissions trade-off index, popularity index, and seasonality index can be revised as follows:
(7) $\mathit{Z(c_{i})}=0.352\cdot\tau_{TT}({\mathit{c_{i}}})+0.218\cdot\tau_{EM}({\mathit{c_{i}}})+0.431\cdot\tau_{Cost}({\mathit{c_{i}}})$
(8) $\rho(c_{i})=0.469\cdot\pi_{POI}+0.325\cdot\pi_{UGC}+0.206\cdot\pi_{GT}$
(9) $\sigma(c_{i}^{j})=0.443\cdot G_{AVC}^{(i)}+0.557\cdot G_{ADR}^{(i)(j)}$
@@
(10) $SFI(c_{i}^{j})=0.281\cdot\mathit{Z(c_{i})}+0.334\cdot\rho(c_{i})+0.385\cdot\sigma(c_{i}^{j})$
@@
Figure 8. Distribution of the absolute values of weights, obtained through the weighted average of Likert scale results
7.3. Integrating S-Fairness Indicator into TRS
@@
Figure 9. Snapshot of the sample user interface showing top travel recommendations from Munich for July while highlighting the S-Fairness Indicator in the top right corner, shown to the participants during the user study
Following the presentation, participants were tasked with expressing their opinions on the provided statements using a 5-point Likert scale (Joshi et al., 2015) ranging from ”strongly disagree” to ”strongly agree”. The statements covered the following aspects:
@@ 3. $S_{3}$:
The results are depicted in Figure 10, revealing a generally positive reception to the S-Fairness Indicator as a sustainability indicator. A majority of users (72%) expressed neutrality or agreement, with 4% of the users expressing very strong agreement. The ”lower value, higher appeal” metric showed a similar trend, with 84% ranging from neutral to solid agreement, including strong agreement, suggesting that a lower value of S-Fairness Indicator correlates with higher user appeal. However, the ”overall helpful metric” elicited more varied responses, with 30% of users agreeing and 32% disagreeing. Given the marginal difference between agreement and disagreement, particularly concerning this metric, it is possible that users did not fully grasp the concept presented in the interface. Improving the representation or providing clearer motivation in the user interface may address this issue in future iterations. This limitation of our study will be addressed in subsequent research efforts. Overall, there is a trend of approval across all metrics, with even the least favorable response showing a majority of users being neutral to strongly agreeing on the value of the metrics.
Figure 10. Results of the user study to validate the usefulness of the S-Fairness Indicator (SFI), showing a mostly positive response

### v3

@@ 5.2 Estimating Seasonal Demand
In the context of seasonal tourism demand, studies suggest that the average room price is one of the pivotal business performance indicators in the hotel industry [41, 56, 62]. We calculate the seasonality Gini index for a city using Gini coefficients derived from non-financial (NFI) and financial indicators (FI), as described by Suštar and Ažić [73]. Non-financial indicators consist of monthly counts of arriving visitors (AVC) from TourMIS, while financial indicators involve the ADR computed from Airbnb listings. The Gini coefficient values for each indicator span from zero to one, with zero signifying a complete lack of seasonality and seems to be an active or equal distribution of volumes all year round. A Gini coefficient of one indicates complete seasonality, i.e., the total volume is registered in one single month [73]. Gini coefficients are computed annually for AVC, while monthly calculations are performed for ADR.
@@
$\sigma(c_{i}^{j})=\Gamma_{AVC}\cdot G_{AVC}^{(i)}+\Gamma_{ADR}\cdot G_{ADR}^{(i)(j)}$ (5)
@@
Table 4: Analysis of the relationship between Airbnb ADR and TourMIS AVC: Pearson’s correlation coefficients and T-Test results for selected European cities
@@ Results
We calculate the Gini coefficients for the AVC and ADR, presenting the results in Table 5. Madrid exhibits the least seasonality in AVC, suggesting consistent demand throughout the year, while Munich registers the highest seasonality. Additionally, Brussels displays minimal seasonality in ADR (close to 0), and Munich exhibits maximum ADR seasonality in September and October. The heightened seasonality in Munich’s AVC and ADR can also be attributed to the annual Oktoberfest event occurring in September [36].
Table 5: Gini coefficients illustrating the seasonality indicators in selected cities. The bold notations highlight the maximum Gini indices for ADR in the city, while the italicized and underlined ones denote the minimum
@@ 5.3 Summary
According to research, the most effective way to measure tourism seasonality is by computing the Gini coefficients of various financial and non-financial indicators [73]. We collected data from TourMIS and Airbnb to analyze non-financial indicators such as monthly arrival visitor counts (AVC) and financial indicators like average daily rates (ADR), respectively. The seasonal demand index is a weighted combination of Gini coefficients calculated from non-financial and financial indicators.
@@
6 User Perception of Sustainable City Trips
To explore how users perceive sustainability when looking for city trip recommendations, we conducted a user study involving participants with diverse travel experiences and preferences. This method provides valuable insights into decision-making intricacies, as established by prior research [81]. By directly engaging real-world participants in simulated scenarios, we aimed to discern how tourists assess different aspects of a city trip, negotiate trade-offs, align preferences with sustainable tourism practices, and assign weights to criteria in their decision-making.
6.1 User Study Design
@@
The questionnaire was designed using Qualtrics Experience Management Software  [^13] https://www.qualtrics.com, an online survey platform. We recruited participants through the online crowdsourcing platform Prolific  [^14] https://www.prolific.com, renowned for its efficacy in subject recruitment for the scientific community [58]. With a focus on European participants who listed travel as one of their hobbies, the questionnaire, designed in English, was distributed to individuals through Prolific’s advanced pre-screening options, and 200 final responses were collected. To ensure gender diversity, the preset distribution aimed for an equal representation of 50% males and 50% females. Demographic analyses of the survey data indicated that 33.8% of the participants fell within the 25-34 age group, followed by 24.3% in the 18-24 age group, and the remaining were above the age of 35.
6.2 Transportation Sustainability Concerns
@@
Figure 4: CO2e, travel time (TT), estimated distance, and cost as shown to the users for each trip scenario, separated by transportation mode. Each city in every trip scenario is denoted by its respective IATA code
Figure 5: Distribution of user responses (%) for different transport modes in various trip scenarios. Each city in every trip scenario is denoted by its respective IATA code
@@
Figure 6: Reasonings provided by respondents when choosing modes of transport across all trips
We investigated the reasons behind selecting trade-offs for various transportation modes in each trip scenario, as illustrated in Figure 6. Our analysis reveals valuable insights from survey responses. For train travel, a significant 47.1% of respondents prioritize convenience, while 15.65% emphasize selecting the mode with the least CO2e emissions. Affordability is a critical factor for 13.91% of participants, and 8.70% opt for the cheapest option available. Additionally, 5.22% favor trains over flights for cost considerations. In contrast, driving is primarily chosen for its convenience and flexibility, with an overwhelming 73.81% of respondents highlighting this aspect. Affordability remains a factor, as 18.57% opt for the cheapest driving option. Some respondents (6.67%) perceive driving as environmentally better than flying. Flying is chosen by 76.53% for its speed, while 11.91% remains unconcerned about CO2e emissions. The findings underline the multifaceted nature of decision-making, encompassing convenience, environmental concerns, and cost considerations across different transportation modes. Our results are consistent with those presented by Avogadro et al [3], which also indicates a preference for public transport over flying for shorter distances.
6.3 Understanding Trade-offs
The user study also aimed to explore the various trade-offs associated with trip planning. Participants were presented with Likert scale [43] statements from "not at all important" to "extremely important" to gauge their agreement levels to the following statements:
@@ 7. $S_{7}$:
Figure 7: Degree of the importance of the various factors influencing decision-making during trip planning
The analysis of user responses, as depicted in Figure 7, provides insights into the trade-offs users are willing to make when considering various attributes in travel decision-making. Notably, for the attribute "popular but crowded," a minimal percentage (2.99%) strongly agreed, while the majority (31.34%) agreed. Conversely, for the attribute "attractive but popular", a significant proportion (55.22%) agreed, indicating a higher tolerance for popularity in the pursuit of attractiveness. The consideration of "overall attractiveness" saw 51.24% in agreement. Respondents expressed varying opinions on the "visiting in peak season", with 1.99% strongly disagreeing and 33.83% expressing disagreement. Furthermore, factors such as "off-season cost savings", "Climate at the destination" during the time of travel, and "off-season discounts" revealed nuanced preferences, with notable percentages in agreement (51.74%, 37.31%, 39.30%, respectively) and distinct proportions holding dissenting views. These results contribute valuable insights into the varied considerations influencing users’ travel preferences.
7 Societal Fairness
This section explores the concept of Societal Fairness (S-Fairness), which is a crucial component in evaluating city trip recommendations. S-Fairness seeks to ensure that tourism benefits and impacts are distributed fairly among a broad range of stakeholders, including not only tourists, service providers, and platforms, but also non-participating entities such as residents, locals, and the environment [6, 8].
@@ 7 Societal Fairness
7.1 Defining the S-Fairness Indicator
@@
$SFI(c_{i}^{j})=\alpha\cdot Z(c_{i})+\beta\cdot\rho(c_{i})+\Gamma\cdot\sigma(c_{i}^{j})$ (6)
@@
7.2 Learning Weights of the Indices
One of the main goals of the user study described in Section 6.1 was to understand the importance of different factors that the users consider when choosing travel destinations. Participants were requested to prioritize their transportation choices based on travel time, CO2e footprint, and cost. They also provided feedback on other essential factors, such as the number of points of interest, Google image search values, and the total number of Tripadvisor reviews and opinions. Additionally, they indicated the importance of costs associated with available accommodation and crowd levels in the city during their visit. These factors align with the coefficients of the "emissions trade-off index" $\mathit{Z(c_{i})}$ in Equation 1, "popularity index" $\rho(c_{i})$ in Equation 3, and "seasonality index" $\sigma(c_{i}^{j})$ in Equation 5 respectively.
@@ 7.2 Learning Weights of the Indices
Figure 8: Distribution of the absolute values of weights, obtained through the weighted average of Likert scale results
We calculated weighted averages on Likert scales, spanning from "not important at all", having a minimum weight of 1, to "extremely important", with a maximum weight of 5, enabling us to identify patterns in the composite indicator. The distribution of the absolute values of these weights, obtained through the weighted average of Likert scale results, are shown in  Figure 8. Additionally, we applied min-max normalization [61] to normalize these averages within each category, enabling us to gauge their relative significance in their respective categories. These normalized weights offer valuable insights into participants’ preferences and priorities, shedding light on the factors that significantly influence their decision-making when choosing travel destinations. After incorporating the normalized weights, the expressions for the emissions trade-off index, popularity index, and seasonality index can be revised as follows:
$\mathit{Z(c_{i})}=0.352\cdot\tau_{TT}({\mathit{c_{i}}})+0.218\cdot\tau_{EM}({\mathit{c_{i}}})+0.431\cdot\tau_{Cost}({\mathit{c_{i}}})$ (7)
$\rho(c_{i})=0.469\cdot\pi_{POI}+0.325\cdot\pi_{UGC}+0.206\cdot\pi_{GT}$ (8)
$\sigma(c_{i}^{j})=0.443\cdot G_{AVC}^{(i)}+0.557\cdot G_{ADR}^{(i)(j)}$ (9)
@@
$SFI(c_{i}^{j})=0.281\cdot\mathit{Z(c_{i})}+0.334\cdot\rho(c_{i})+0.385\cdot\sigma(c_{i}^{j})$ (10)
@@
7.3 Integrating S-Fairness Indicator into TRS
@@
Figure 9: Snapshot of the sample user interface showing top travel recommendations from Munich for July while highlighting the S-Fairness Indicator in the top right corner, shown to the participants during the user study
Following the presentation, participants were tasked with expressing their opinions on the provided statements using a 5-point Likert scale [43] ranging from "strongly disagree" to "strongly agree". The statements covered the following aspects:
@@ 3. $S_{3}$:
The results are depicted in Figure 10, revealing a generally positive reception to the S-Fairness Indicator as a sustainability indicator. A majority of users (72%) expressed neutrality or agreement, with 4% of the users expressing very strong agreement. The "lower value, higher appeal" metric showed a similar trend, with 84% ranging from neutral to solid agreement, including strong agreement, suggesting that a lower value of S-Fairness Indicator correlates with higher user appeal. However, the "overall helpful metric" elicited more varied responses, with 30% of users agreeing and 32% disagreeing. Given the marginal difference between agreement and disagreement, particularly concerning this metric, it is possible that users did not fully grasp the concept presented in the interface. Improving the representation or providing clearer motivation in the user interface may address this issue in future iterations. This limitation of our study will be addressed in subsequent research efforts. Overall, there is a trend of approval across all metrics, with even the least favorable response showing a majority of users being neutral to strongly agreeing on the value of the metrics.
Figure 10: Results of the user study to validate the usefulness of the S-Fairness Indicator (SFI), showing a mostly positive response